Christopher’s Custom Cabinet Company uses a job order cost system with overhead applied as a percentage of direct labor costs. Inventory balances at the beginning of 2018 follow:

Raw Materials Inventory $ 16,300
Work in Process Inventory 5,100
Finished Goods Inventory 21,600


The following transactions occurred during January:


(a) Purchased materials on account for $27,800.
(b) Issued materials to production totaling $21,700, 90 percent of which was traced to specific jobs and the remainder of which was treated as indirect materials.
(c) Payroll costs totaling $17,800 were recorded as follows:
    $11,000 for assembly workers
     2,900 for factory supervision
     1,700 for administrative personnel
      2,200 for sales commissions
(d) Recorded depreciation: $5,400 for factory machines, $600 for the copier used in the administrative office.
(e) Recorded $1,000 of expired insurance. Forty percent was insurance on the manufacturing facility, with the remainder classified as an administrative expense.
(f) Paid $5,900 in other factory costs in cash.
(g) Applied manufacturing overhead at a rate of 200 percent of direct labor cost.
(h) Completed all jobs but one; the job cost sheet for the uncompleted job shows $2,300 for direct materials, $2,100 for direct labor, and $4,200 for applied overhead.
(i) Sold jobs costing $50,900. The revenue earned on these jobs was $66,170.

Required:
1. Set up T-accounts, record the beginning balances, post the January transactions, and compute the final balance for the following accounts:

  1. Raw Materials Inventory.
  2. Work in Process Inventory.
  3. Finished Goods Inventory.
  4. Cost of Goods Sold.
  5. Manufacturing Overhead.
  6. Selling, General, and Administrative Expenses.
  7. Sales Revenue.

2. Determine how much gross profit the company would report during the month of January before any adjustment is made for the overhead balance.

3. Determine the amount of over- or underapplied overhead.

4. Compute adjusted gross profit assuming that any over- or underapplied overhead balance is adjusted directly to Cost of Goods Sold.Raw Materials Inventory 16.300 Work in Process Inventory 5.100 Beg. Bal. Beg Bal End. Bal. End. Bal. Cost of Goods Sold Finis
Answer #1

Answer 1:

Raw Material Inventory $16,300 $21,700 $27,800 (b Beg. Bal. __ (a) (h) Work in Process Inventory Beg. Bal. $5,100 $49,030 $19

Calculations:

(b)

Direct material = 21700 * 90% = $19,530

Indirect material = 21700 - 19530 = $2170

(e)

Insurance:

Manufacturing overhead = 1000 * 40% = $400

Selling, General and Administrative Overhead = 1000 - 400 = $600

(g) Applied manufacturing overhead = 200% *  direct labor cost = 200% * 11000 = $22,000

(h) Ending balance of work in process = 2300 + 2100 + 4200 = $8,600

Finished goods transferred from work in process = Balancing figures = 5100 + 19530 + 11000 + 22000 - 8600 = $49,030

Answer 2:

Unadjusted Gross profit $15,270

Working:

Unadjusted Gross profit = Sales - Cost of unadjusted goods sold = 66170 -50900 = $15,270

Answer 3:

Manufacturing overhead $5,230 over-applied

Working:

From T Account of Manufacturing Overhead, we find credit balance of $5230 which is over-applied overhead

Answer 4:

Adjusted Gross profit $20,500

Working:

Adjusted cost of goods sold = 50900 - 5230 =$45,670

Adjusted gross profit = Sales - Adjusted cost of goods sold = 66170 - 45670 = $20,500